OpenStudy (anonymous):

Solve for r, s, and t 6r-s+t=31 4r+2-3t=34 r-3s+2t=-9

OpenStudy (anonymous):

6r-s+t=31 solve for t in this first one: add -6r +s to both sides to get: t= 31-6r +s insert 31-6r +s in place of t to solve for another variable place that value in to get exact value of last variable, substitute and find exact valur of remaining.

OpenStudy (anonymous):

thanks